2 runners are running a track one runner runs it in 24 min, the other runner in 36 minutes when will they both be at the same spot?
Found 2 solutions by ankor@dixie-net.com, checkley77:
Answer by ankor@dixie-net.com(22740)

2 runners are running a track one runner runs it in 24 min,
the other runner in 36 minutes when will they both be at the same spot?
:
Prime factor both values
24: 2*2*2*3
36: 2*2*3*3
LCM 2*2*2*3*3 = 72 minutes
:
one runner completes 3 laps
one runner completes 2 laps

Answer by checkley77(12844)

Find the LCD for 24 & 36
12*2=24
12*3=36
12*2*3=72 is the LCM.
Thus in 72 minutes they will meet.